# A Survey on Resource Management in Joint Communication and Computing-Embedded SAGIN

### v1

@@ I-A Background and Motivation
Terrestrial cellular networks, boasting high speeds, low latency, and extensive connectivity, can fulfill the communication requirements in typical urban and suburban locales. In the present information era, the communication scope has expanded to remote regions, including oceans, volcanoes, deserts, and forests. Nonetheless, the establishment of TN in these areas is impeded by commercial and operational considerations. Presently, terrestrial cellular networks can cover a mere 7% of the Earth’s surface, leaving an estimated 3.4 billion individuals without network access. This falls significantly short of the objective of seamless global coverage. Hence, 6G envisions the incorporation of NTN, utilizing satellites, unmanned aerial vehicles (UAVs), and high altitude platforms (HAPs) to supplement terrestrial coverage. The construction of multidimensional SAGIN is crucial to achieve comprehensive area coverage. For the sake of brevity, this paper refers to non-terrestrial platforms as space-air platforms (SAPs).
@@
With the implementation of onboard processing in SAPs, the possibility of integrating communication and computing within NTN becomes both feasible and promising. The architecture of a JCC-SAGIN is illustrated in Fig. 2. This approach is particularly advantageous for latency-sensitive or computation-intensive services, especially in remote areas. Such services, traditionally limited by the computational capacities of terrestrial IoT devices or reliant on remote cloud computing at ground stations (GSs), can now be offloaded to visible SAPs for edge computing. JCC-SAGIN, representing a significant advancement in the field, necessitates a comprehensive review and summarization of its research issues and challenges to fully realize its potential.
@@ I-B1 Surveys on NTN
Ref. Year Architecture SAP as Edge Computing Server Contribution
[1] 2018 SAGIN $\times$ The first extensive overview of the integration of space, air, and ground networks, delving into the research challenges, technologies, and applications in this emerging field.
[2] 2020 Satellite $\times$ A literature review encompassing a range of CubeSat missions and recent developments in constellation-and-coverage issues, channel modeling, physical techniques, and networking.
[3] 2020 Satellite $\times$ A thorough survey that focuses on space information networks, particularly highlighting PLS.
@@
[11] 2022 UAV $\times$ A survey that examines UAV communications from an industrial viewpoint, analyzing the potential and limitations of 5G NR features in aerial devices and identifying promising 6G enablers for UAV communication.
[12] 2022 UAV $\times$ A detailed analysis focusing on mmWave beamforming-enabled UAV communications and networking.
[13] 2022 SAGSIN $\times$ The first survey to critically review state-of-the-art security in SAGSIN, including discussions on cross-layer attacks and security countermeasures.
[14] 2023 NTN $\times$ A comprehensive survey reviewing both single-tier and multi-tier scenarios in aerospace integrated networks.
This paper 2024 SAGIN An all-encompassing overview that discusses the network and computing architecture, resource management strategies, security issues, and future research directions in JCC within SAGIN.
@@ I-B2 Surveys on the Integration of Communication with Other Technologies
Recent advancements in integrated sensing and communication (ISAC) have highlighted its potential to optimize the use of spatial, temporal, frequency, and power resources. This optimization comes alongside reductions in hardware and software complexity within wireless communications and radar sensing domains. The core of ISAC’s effectiveness lies in its signal processing capabilities, which are pivotal for enhancing both sensing and communication functions. Wei et al. [16] conducted a thorough examination of ISAC signal methodologies in the context of 5G-Advanced and 6G networks, addressing signal design, processing, and optimization, alongside methods for radar signal processing and interference management strategies. Utilizing the distinct benefits of MIMO technology, Fang et al. [17] explored advancements in ISAC via spatial beamforming and waveform shaping, incorporating novel MIMO models with cloud radio access networks, UAVs, and reconfigurable intelligent surfaces (RISs). Luong et al. [18] presented an extensive survey on the resource management challenges inherent in ISAC systems. Concurrently, Ma et al. [19] reviewed the integration of IoT sensing, computing, and communication, with a particular focus on advancements in energy harvesting (EH) for IoT systems.
@@
Ref. Year Network Type Technique Contribution
@@
[17] 2023 IoT Sensing and communication A survey specializing in ISAC applications using multiple-input multiple-output (MIMO) technology. This review discusses fundamental models combined with cloud-radio access networks (C-RANs), UAVs, and reflective intelligent surfaces (RISs), highlighting the potential and challenges of ISAC MIMO designs.
@@
This paper 2024 SAGIN Networking, communication and computing A survey on the theoretical and engineering advancements in network architecture, application areas, and optimization methods within the context of JCC-SAGIN.
@@
  • Scalability and Integration: NTN must accommodate extensive geographic coverage and a large user base, presenting scalability challenges in routing and network management. They also require seamless integration with TN and other communication systems, alongside the development of interoperable standards, protocols, and architectures.
@@ I-B3 Surveys on Resource Management
Ref. Year Network Type Hardware Applications Traditional Alg. AI Alg. Contribution
[24] 2022 5G HetNets $\times$ $\times$ A discussion on the critical role of RRM in enhancing 5G HetNets through addressing challenges such as interference and resource allocation, presenting a survey of RRM schemes, and outlining future challenges and opportunities.
[26] 2022 MEC system $\times$ $\times$ $\times$ An overview of machine learning and deep learning applications for resource allocation in MEC systems, detailing their use in task offloading, scheduling, and joint resource allocation, and discussing future challenges and directions in this area.
[28] 2021 Big data networks $\times$ $\times$ A comprehensive review of big data deployment architectures, introducing a taxonomy for classifying these models based on their communication systems, as well as the implications, benefits, and challenges of modern big data environments compared to traditional dedicated clusters.
[23] 2020 5G and Beyond $\times$ $\times$ $\times$ A survey paper reviews research and standardization efforts aimed at addressing cross-link interference in dynamic time division duplex (D-TDD) systems for 5G New Radio, categorizing mitigation approaches, discussing signaling requirements, and presenting performance analysis in various environments.
[25] 2020 Cellular and IoT Networks $\times$ $\times$ An in-depth survey of machine learning and deep learning based resource management mechanisms in cellular wireless and IoT networks, addressing the challenges of resource management, reviewing traditional and ML/DL-based techniques, and identifying future research directions for enhancing IoT networks’ efficiency and intelligence.
[27] 2020 MEC system $\times$ $\times$ A survey presents the application of machine learning algorithms in MEC systems to address high-dimensional challenges in task offloading, resource allocation, and inter-server communication for 5G and IoT networks.
This paper 2024 SAGIN We provide a comprehensive survey of computing hardware, JCC applications, traditional and AI algorithms of resource management optimization in JCC-SAGIN.
@@ I-C Organization of This Survey
Fig. 4 illustrates the structural organization of this survey. Following the introductory section, which illustrates the context and reviews pertinent existing surveys, a thorough examination of JCC-SAGIN is undertaken from Section II to Section VII. Specifically, Section II summarizes the architecture and evolution of JCC-SAGIN, tracing the network components and the evolution of network integration and computing hardware in JCC-SAGIN. Section III discusses the enabling technologies, encompassing network slicing, software-defined networking (SDN), network function virtualization (NFV), EH and wireless power transfer (WPT), and RIS within JCC-SAGIN. Building upon the foundational architecture and promising techniques, Section IV explores the typical applications facilitated by JCC-SAGIN and highlights engineering cases that have been successfully implemented. The survey progresses from Section V to Section VI, where we cover resource management modeling and optimization methods in JCC-SAGIN. In Section VII, we identify future research opportunities, paving the way for further exploration and innovation in this field. Our work is summarized in Section VIII.
@@ II-A1 Space Segment
The space segment of SAGIN primarily comprises satellites. The satellites are categorized based on their orbital heights into three types: geostationary (GEO), medium earth orbit (MEO), and low earth orbit (LEO). Table IV delineates the distinct characteristics of satellites operating at varying orbital altitudes, enabling the formation of hierarchical multi-layer satellite networks.
TABLE IV: Comparison of GEO, MEO, and LEO satellites
@@
Data rate High Not as high as GEO. High, often match to or even exceed those of GEO satellites.
Advantages Global visibility. Fixed coverage area. Long-term stable communication. Relatively lower cost. Cost-effective, rapid deployment. Suitable for large-scale constellations.
Disadvantages High cost and complexity. Limited bandwidth. Large Doppler shift. Possible interruptions in communication.
@@
TABLE V: Typical LEO satellite constellations
@@
Table V provides a comprehensive summary of notable LEO satellite constellations. A significant evolution in satellite functionality is observed, transitioning from transparent forwarding to onboard processing capabilities. For instance, Starlink Internet project of SpaceX involves the deployment of 32,000 Linux computers and over 6,000 microcontrollers in the space[30]. This advancement underscores the necessity to consider the coexistence of satellites with varying functionalities within the space segment, particularly when developing communication strategies. Moreover, the comparatively shorter distance between LEO satellites and mobile terminals as opposed to GEO and MEO satellites facilitates direct communication. This has spurred various satellite constellation programs aimed at providing global service coverage directly to mobile phones. Key developments in this area are listed as follows:
@@ II-A2 Air Segment
Furthermore, the tens of thousands of CAs traversing the sky offer a type of promising aerial platforms. Since CAs operate on existing flight routes, they do not incur additional launching costs. In comparison to UAVs, CAs face fewer constraints regarding load and energy capacity, and a single CA can provide substantial coverage. Modern developments enable passengers to access internet services on board through air-to-ground (ATG) connectivity, facilitated by communication links with LEO satellites or ground gateways. Several commercial and trial networks are operational, such as Gogo’s network in the USA, Inmarsat’s in Europe, and CMCC’s trial network in China, as referenced in [33]. Standards pertinent to these networks are poised for future implementation.Nonetheless, the flight trajectories of CAs are inherently less flexible than UAVs, being dependent on predefined flight routes. Additionally, CA trajectories are less stable compared to satellites, as they can be influenced by environmental factors and other variables.
@@ II-B Evolution of Computing Hardware in JCC-SAGIN
II-B1 High-Performance Spaceflight Computing (HPSC) Processor
In the aerospace sector, chip reliability is paramount, even more so than performance. Only chips with radiation resistance can guarantee the normal functioning of spacecraft. Numerous suppliers are producing highly integrated and modular onboard computing systems tailored for small spacecraft. Table VI outlines some of the latest advancements in aerospace-grade highly integrated airborne computing products [37]. The vehicle column in the table correlates each onboard unit with a specific spacecraft classification, where “general satellite” refers to larger SmallSat platforms, exceeding the size of CubeSats. The technology readiness level of a chip indicates its maturity for space applications, which can vary based on specific payload and mission requirements.
TABLE VI: Sample of Highly Integrated Onboard Computing Systems
@@ II-B1 High-Performance Spaceflight Computing (HPSC) Processor
With the advent of HPSC processors, the onboard processing capabilities of satellite networks have seen significant enhancement. Table VII provides a comprehensive summary of typical LEO satellite networks equipped with onboard computing capabilities, supported by advanced aerospace-grade entral processing unit (CPU) chips [38]. Fig. 5 presents a comparative analysis of the computing power of chips utilized for onboard data processing. The metrics used for this comparison are Million Instructions Per Second (MIPS) and Giga Instructions Per Second (GIPS), which measure the computing power of a CPU by indicating the number of instructions it can execute per second. This comparison provides valuable insights into the processing capabilities of various chips, highlighting their suitability for onboard data processing tasks in satellite networks.
TABLE VII: Typical LEO Satellite Networks with onboard processing
Satellite System Year Country Satellite Networking Onboard Processing Function
@@
Tianzhi Satellite 2019 China $\times$ Cloud detection and satellite control
Tianxian Constellation 2021 $\times$ On-orbit processing of synthetic-aperture radar (SAR) images
Tiansuan Satellite 2021 $\times$ Target recognition and model training

### v3

@@ I-A Background and Motivation
Terrestrial cellular networks, boasting high speeds, low latency, and extensive connectivity, can fulfill the communication requirements in typical urban and suburban locales. In the present information era, the communication scope has expanded to remote regions, including oceans, volcanoes, deserts, and forests. Nonetheless, the establishment of terrestrial networks (TN) in these areas is impeded by commercial and operational considerations. Presently, TN can cover a mere 7% of the Earth’s surface, leaving an estimated 3.4 billion individuals without network access. This falls significantly short of the objective of seamless global coverage. Hence, 6G envisions the incorporation of non-terrestrial networks (NTN), utilizing satellites, unmanned aerial vehicles (UAVs), and high altitude platforms (HAPs) to supplement terrestrial coverage. The construction of multidimensional space-air-ground integrated networks (SAGIN) is crucial to achieve comprehensive area coverage. For the sake of brevity, this paper refers to non-terrestrial platforms as space-air platforms (SAPs).
@@
With the implementation of onboard processing in SAPs, the possibility of integrating communication and computing within NTN becomes both feasible and promising. The architecture of joint communication and computing-embedded SAGIN (JCC-SAGIN) is illustrated in Fig. 2. This approach is particularly advantageous for latency-sensitive or computation-intensive services, especially in remote areas. Such services, traditionally limited by the computational capacities of terrestrial IoT devices or reliant on remote cloud computing at ground stations (GSs), can now be offloaded to visible SAPs for edge computing. JCC-SAGIN, representing a significant advancement in the field, necessitates a comprehensive review and summarization of its research issues and challenges to fully realize its potential.
@@ I-B1 Surveys on NTN
Ref. Year Network SAP as Edge Computing Server Contribution
[1] 2018 SAGIN The first extensive overview of the integration of space, air, and ground networks, delving into the research challenges, technologies, and applications in this emerging field.
[2] 2020 Satellite A literature review encompassing a range of CubeSat missions and recent developments in constellation-and-coverage issues, channel modeling, physical techniques, and networking.
[3] 2020 Satellite A thorough survey that focuses on space information networks, particularly highlighting PLS.
@@
[11] 2022 UAV A survey that examines UAV communications from an industrial viewpoint, analyzing the potential and limitations of 5G NR features in aerial devices and identifying promising 6G enablers for UAV communication.
[12] 2022 UAV A detailed analysis focusing on mmWave beamforming-enabled UAV communications and networking.
[13] 2022 SAGSIN The first survey to critically review state-of-the-art security in SAGSIN, including discussions on cross-layer attacks and security countermeasures.
[14] 2023 NTN A comprehensive survey reviewing both single-tier and multi-tier scenarios in aerospace integrated networks.
Ours 2024 SAGIN An all-encompassing overview that discusses the network and computing architecture, resource management strategies, security issues, and future research directions in JCC within SAGIN.
@@ I-B2 Surveys on the Integration of Communication with Other Technologies
Recent advancements in integrated sensing and communication (ISAC) have highlighted its potential to optimize the use of spatial, temporal, frequency, and power resources. This optimization comes alongside reductions in hardware and software complexity within wireless communications and radar sensing domains. The core of ISAC’s effectiveness lies in its signal processing capabilities, which are pivotal for enhancing both sensing and communication functions. Wei et al. [16] conducted a thorough examination of ISAC signal methodologies in the context of 5G-Advanced and 6G networks, addressing signal design, processing, and optimization, alongside methods for radar signal processing and interference management strategies. Utilizing the distinct benefits of multiple-input multiple-output (MIMO) technology, Fang et al. [17] explored advancements in ISAC via spatial beamforming and waveform shaping, incorporating novel MIMO models with cloud radio access networks, UAVs, and reconfigurable intelligent surfaces (RISs). Luong et al. [18] presented an extensive survey on the resource management challenges inherent in ISAC systems. Concurrently, Ma et al. [19] reviewed the integration of IoT sensing, computing, and communication, with a particular focus on advancements in energy harvesting (EH) for IoT systems.
@@
Ref. Year Network Technique Contribution
@@
[17] 2023 IoT Sensing and communication A survey specializing in ISAC applications using MIMO technology. This review discusses fundamental models combined with cloud-radio access networks (C-RANs), UAVs, and RISs, highlighting the potential and challenges of ISAC MIMO designs.
@@
Ours 2024 SAGIN Networking, communication and computing A survey on the theoretical and engineering advancements in network architecture, application areas, and optimization methods within the context of JCC-SAGIN.
@@
  • Scalability and Integration: NTN must accommodate extensive geographic coverage and a large user base, presenting scalability challenges in routing and network management. They also require seamless integration with other communication systems, alongside the development of interoperable standards, protocols, and architectures.
@@ I-B3 Surveys on Resource Management
Ref. Year Network Hardware Applications Traditional Alg. AI Alg. Contribution
[24] 2022 5G HetNets A discussion on the critical role of RRM in enhancing 5G HetNets through addressing challenges such as interference and resource allocation, presenting a survey of RRM schemes, and outlining future challenges and opportunities.
[26] 2022 MEC system An overview of machine learning and deep learning applications for resource allocation in MEC systems, detailing their use in task offloading, scheduling, and joint resource allocation, and discussing future challenges and directions in this area.
[28] 2021 Big data networks A comprehensive review of big data deployment architectures, introducing a taxonomy for classifying these models based on their communication systems, as well as the implications, benefits, and challenges of modern big data environments compared to traditional dedicated clusters.
[23] 2020 5G and beyond A survey paper reviews research and standardization efforts aimed at addressing cross-link interference in dynamic time division duplex (D-TDD) systems for 5G New Radio, categorizing mitigation approaches, discussing signaling requirements, and presenting performance analysis in various environments.
[25] 2020 Cellular and IoT networks An in-depth survey of machine learning and deep learning based resource management mechanisms in cellular wireless and IoT networks, addressing the challenges of resource management, reviewing traditional and learning-based techniques, and identifying future research directions for enhancing IoT networks’ efficiency and intelligence.
[27] 2020 MEC system A survey presents the application of machine learning algorithms in MEC systems to address high-dimensional challenges in task offloading, resource allocation, and inter-server communication for 5G and IoT networks.
Ours 2024 SAGIN We provide a comprehensive survey of computing hardware, JCC applications, traditional and AI algorithms of resource management optimization in JCC-SAGIN.
@@ I-C Organization of This Survey
For readability, Table IV summarizes the primary acronyms. Fig. 4 illustrates the structural organization of this survey. Following the introductory section, which illustrates the context and reviews pertinent existing surveys, a thorough examination of JCC-SAGIN is undertaken from Section II to Section VII. Specifically, Section II summarizes the architecture and evolution of JCC-SAGIN, tracing the network components and the evolution of network integration and computing hardware in JCC-SAGIN. Section III discusses the enabling technologies, encompassing network slicing, software-defined networking (SDN), network function virtualization (NFV), EH and wireless power transfer (WPT), and RIS within JCC-SAGIN. Building upon the foundational architecture and promising techniques, Section IV explores the typical applications facilitated by JCC-SAGIN and highlights engineering cases that have been successfully implemented. The survey progresses from Section V to Section VI, where we cover resource management modeling and optimization methods in JCC-SAGIN. In Section VII, we identify future research opportunities, paving the way for further exploration and innovation in this field. Our work is summarized in Section VIII.
TABLE IV: LIST OF MAIN ABBREVIATIONs
Acronyms Description Acronyms Description
AC Active Communication MATD3 Multi-Agent TD3
AI Artificial Intelligence MDP Markov Decision Process
AP Access Point MEC Mobile Edge Computing
BCD Block Coordinate Descending MEO Medium Earth Orbit
BPP Binomial Point Process MINLP Mixed Integer Nonlinear Programming
BS Base Station ML Machine Learning
CA Civil Airplane NFV Network Function Virtualization
CDC Coded Distributed Computing NPPP Nonhomogeneous Poisson Point Process
CSI Channel State Information NTN Non-Terrestrial Networks
CTCE Centralized Training and Centralized Execution PCP Poisson Cluster Process
CTDE Centralized Training and Decentralized Execution PLS Physical Layer Security
D3QN Dueling Double DQN PPP Poisson Point Process
DC Difference of Convex QKD Quantum Key Distribution
DDNN Distributed DNN QL Q-Learning
DDPG Deep Deterministic Policy Gradient QoE Quality of Experience
DDQN Double DQN QoS Quality of Service
DL Deep Learning QT Quadratic Transform
DNN Deep Neural Network RIS Reconfigurable Intelligent Surface
DQL Deep Q-Learning RL Reinforcement Learning
DQN Deep Q-Network SAGIN Space-Air-Ground Integrated Networks
DRL Deep Reinforcement Learning SAP Space-Air Platform
DT Digital Twin SCA Successive Convex Approximation
DTDE Decentralized Execution SDN Software-Defined Networking
EE Energy Efficiency SDR Semi-Definite Relaxation
EH Enengy Harvesting SR Shadowed-Rician
FL Federated Learning STIN Satellite-Terrestrial Integration Networks
FP Fractional Programming TD Temporal Difference
GEO Geostationary TL Transfer Learning
GS Ground Station TN Terrestrial Network
HAP High Altitude Platform UAV Unmanned Aerial Vehicle
IoT Internet of Things URLLC Ultra-Reliable and Low Latency Communications
ISAC Integrated Sensing and Communication VNF Virtual Network Function
JCC-SAGIN Joint Communication and Computing-Embedded SAGIN VR Virtual Reality
LEO Low Earth Orbit VSAT Very Small Aperture Terminal
MADDPG Multi-Agent DDPG WPT Wireless Power Transfer
@@ II-A1 Space Segment
The space segment of SAGIN primarily comprises satellites. The satellites are categorized based on their orbital heights into three types: geostationary (GEO), medium earth orbit (MEO), and low earth orbit (LEO). Table V delineates the distinct characteristics of satellites operating at varying orbital altitudes, enabling the formation of hierarchical multi-layer satellite networks.
TABLE V: Comparison of GEO, MEO, and LEO satellites
@@
Data rate High Not as high as GEO High, often match to or even exceed those of GEO satellites
Advantages Global visibility, fixed coverage area, and long-term stable communication. Relatively lower cost Cost-effective, rapid deployment, and suitable for large-scale constellations.
Disadvantages High cost and complexity Limited bandwidth Large Doppler shift. Possible interruptions in communication.
@@
TABLE VI: Typical LEO satellite constellations
@@
Table VI provides a comprehensive summary of notable LEO satellite constellations. A significant evolution in satellite functionality is observed, transitioning from transparent forwarding to onboard processing capabilities. For instance, Starlink Internet project of SpaceX involves the deployment of 32,000 Linux computers and over 6,000 microcontrollers in the space[30]. This advancement underscores the necessity to consider the coexistence of satellites with varying functionalities within the space segment, particularly when developing communication strategies. Moreover, the comparatively shorter distance between LEO satellites and mobile terminals as opposed to GEO and MEO satellites facilitates direct communication. This has spurred various satellite constellation programs aimed at providing global service coverage directly to mobile phones. Key developments in this area are listed as follows:
@@ II-A2 Air Segment
Furthermore, the tens of thousands of CAs traversing the sky offer a type of promising aerial platforms. Since CAs operate on existing flight routes, they do not incur additional launching costs. In comparison to UAVs, CAs face fewer constraints regarding load and energy capacity, and a single CA can provide substantial coverage. Modern developments enable passengers to access internet services on board through air-to-ground connectivity, facilitated by communication links with LEO satellites or ground gateways. Several commercial and trial networks are operational, such as Gogo’s network in the USA, Inmarsat’s in Europe, and CMCC’s trial network in China, as referenced in [33]. Standards pertinent to these networks are poised for future implementation.Nonetheless, the flight trajectories of CAs are inherently less flexible than UAVs, being dependent on predefined flight routes. Additionally, CA trajectories are less stable compared to satellites, as they can be influenced by environmental factors and other variables.
@@ II-B Evolution of Computing Hardware in JCC-SAGIN
II-B1 High-Performance Spaceflight Computing (HPSC) Processors
In the aerospace sector, chip reliability is paramount, even more so than performance. Only chips with radiation resistance can guarantee the normal functioning of spacecraft. Numerous suppliers are producing highly integrated and modular onboard computing systems tailored for small spacecraft. Table VII outlines some of the latest advancements in aerospace-grade highly integrated airborne computing products [37]. The vehicle column in the table correlates each onboard unit with a specific spacecraft classification, where “general satellite” refers to larger SmallSat platforms, exceeding the size of CubeSats. The technology readiness level of a chip indicates its maturity for space applications, which can vary based on specific payload and mission requirements.
TABLE VII: Sample of Highly Integrated Onboard Computing Systems
@@ II-B1 High-Performance Spaceflight Computing (HPSC) Processors
With the advent of HPSC processors, the onboard processing capabilities of satellite networks have seen significant enhancement. Table VIII provides a comprehensive summary of typical LEO satellite networks equipped with onboard computing capabilities, supported by advanced aerospace-grade entral processing unit (CPU) chips [38]. Fig. 5 presents a comparative analysis of the computing power of chips utilized for onboard data processing. The metrics used for this comparison are million instructions per second (MIPS) and giga instructions per second (GIPS), which measure the computing power of a CPU by indicating the number of instructions it can execute per second. This comparison provides valuable insights into the processing capabilities of various chips, highlighting their suitability for onboard data processing tasks in satellite networks.
TABLE VIII: Typical LEO Satellite Networks with onboard processing
Satellite System Year Country Networking Onboard Processing Function
@@
Tianzhi Satellite 2019 China Cloud detection and satellite control
Tianxian Constellation 2021 On-orbit processing of synthetic-aperture radar (SAR) images
Tiansuan Satellite 2021 Target recognition and model training